OpenStudy (anonymous):

(13)fourthroot of 36 minus (6)sqrt. of 216 plus (3) sixthroot of 216

OpenStudy (anonymous):

\[13\sqrt{\sqrt[4]{36}}-6\sqrt{216}+3\sqrt[6]{216}= ?\]

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=16*sqrt%28sqrt [4]36%29-6*sqrt%28216%29%2B3*%28sqrt[6]216%29 ?